Two sets of data are given. Compute their Sd(variance difference)

X1

42 53 48 65 40 52  

X2

39 45 40 58 42 47



1
Expert's answer
2021-05-07T09:34:58-0400

For X_1

Mean μ  = Σxi/N

μ =42+53+48+65+40+52=300/6

=50

Standard deviation δ = sqrt δ2

. δ2= 64+9+4+225+100+4=406/6

. =67.67

δ= sqrt 67.67=8.23

Standard deviation =8.23

For X_2

Mean μ =(39+45+40+42+47+58)/6=45.2

δ2 = 17.64+0.04+27.04+163.84+10.2

=222.04/6

=37.01

δ= sqrt 37.01

=6.08

Standard deviation is 6.08